# State-witness contraction

### v1

@@ V Multiple copies
VI Distillability
By the Choi-Jamiołkowski isomorphism, each state or witness opertor $R_{j}$ can be seen as the Choi operator of a map $\Psi_{j}(\varrho_{j}^{T}):=\operatorname{tr}_{1...k_{j}-1}(R^{(j)}\cdot{\varrho^{j}}\otimes\mathds{1}_{k})$ from the space of operators acting on $\mathds{C}^{d_{1}^{j}}\otimes...\otimes\mathds{C}^{d_{k_{j}-1}}$ to the space of operators acting on $\mathds{C}^{d_{k_{j}}}$, whose output is positive semidefinite if the inputs are products of positive semidefinite operators (further details are given in [21]). In the bipartite case, consider $R^{(1)}=P_{A}\Pi^{\otimes k}$ and $Q_{B}:=R^{(2)}=Q_{B}\Pi^{\otimes k}$ where $P_{A}$ and $Q_{B}$ are projectors onto a aubspace of $\mathds{C}^{d_{1}^{j}}\otimes...\otimes\mathds{C}^{d_{k_{j}}}$, and $\Pi$ is a local projector onto $\mathds{C}^{2}$ (see Tab. I and Fig. 2 in [21] for specific results derived in this case). Then the action of $\Psi_{A_{1}...A_{k-1}}\otimes\Psi_{B_{1}...B_{k-1}}(\varrho_{AB}^{\otimes k-1})$ projects $k-1$ copies to the two-qubit subspace $\mathds{C}^{2}\otimes\mathds{C}^{2}$, where where the PPT-criterion detects both entanglement and distillability [46, 43].
The framework introduced in this work naturally provides a criterion for $k$-copy distillability, since a state is distillable if and only if it can be transformed into a two-qubit entangled state through local projections on $k$ copies [47, 48]. This shall be illustrated in specific examples involving a single copy, considering the two-qudit antisymmetrizer $P_{1,1}=(\eta_{d}()-\eta_{d}(12))/2$ and a two-dimensional local projector $\Pi$ diagonal in the computational basis: we demonstrate that most generic states in a random sample are detected to be one-copy distillable by using $R^{(1)}=P_{1,1}\Pi^{\otimes 2}$ (see [21] for a detailed numerical analysis), effectively detecting the qubit-qudit distillable entanglement present in a subspace of a two-qudit system. Moreover, choosing $R^{(1)}=R^{(2)}=P_{1,1}\Pi^{\otimes 2}$, the operator $\mathcal{W}=R^{(1)}_{A_{1}A_{2}}\otimes R^{(2)}_{B_{1}B_{2}}$ detects both entanglement and one-copy distillability on all Werner states $\varrho_{W}$ known to be distillable, namely
@@ VI Distillability
within $p\in(p_{0},1/2]$ with $p_{0}=\frac{d+1}{4d-2}$ [49], where $d_{s}=d(d+1)/2$ and $d_{s}=d(d-1)/2$. Indeed, one verifies that
$\operatorname{tr}_{A_{1}B_{1}}\Big{(}\mathcal{W}\cdot{\varrho_{W}}_{A_{1}B_{1}}\otimes\mathds{1}_{A_{2}B_{2}}\Big{)}=pP_{a}^{(2)}/d_{s}+(1-p)P_{s}^{(2)}/d_{a}\,.$ (9)
Therefore, choosing the two-qubit SWAP operator $\tau=2|\phi^{+}\rangle\langle\phi^{+}|^{T_{A}}=\eta_{2}(12)$, for $p\in(p_{0},1/2]$ we have $\operatorname{tr}(\mathcal{W}_{\tau}\varrho_{W})<0$.
VII Conclusions
@@
We have developed a novel technique to design new entanglement witnesses, and demonstrated its capabilities with examples. As a proof of concept we have shown that the method allows to use witnesses unable to detect locally bound entangled states, to construct new larger witnesses able to do so. Moreover, the witnesses constructed in this way can be naturally be optimized to desired states to be detected. This can be done with semidefinite programming whose variable size is smaller than the state to be detected. As a very special case, the method applies to construct nonlinear witnesses acting on multiple copies. With little computational resources, we have seen that our previous multicopy construction provided in [20] can be tuned to significantly better detect noisy entangled states, while the desired properties of unitary invariance and implementability with randomized measurements have been shown to be preserved. Moreover, the technique introduced leads to practical recipes to detect distillability, which are demonstrated to be effective for relevant families of states.
The examples provided demonstrate that this technique opens an avenue to constructing both linear and nonlinear witnesses with desired properties, systematically tailored for a variety of purposes. In particular, the efficient implementability of our method relies on optimizing over a variable $\tau$ restricted to be a state or a witness that is decomposable across some bipartition. For further research, it would be interesting to analyze whether fully non-decomposable witnesses can be obtained when the operator $\tau$ is chosen to be fully non-decomposable; and whether optimal witnesses in the sense of [50] can be systematically obtained.
@@
Acknowledgements:   The author is thankful to Felix Huber, Robin Krebs, Ferran Riera-Sàbat, Some Sankar Bhattacharya and Anna Sanpera for comments and support. Support by the Foundation for Polish Science through TEAM-NET project POIR.04.04.00-00-17C1/18-00 and by NCN QuantERA Project No. 2021/03/Y/ST2/00193 is gratefully acknowledged. The author also acknowledges financial support from Spanish MICIN (projects: PID2022:141283NBI00;139099NBI00) with the support of FEDER funds, the Spanish Goverment with funding from European Union NextGenerationEU (PRTR-C17.I1), the Generalitat de Catalunya, the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project -Quantum Spain Project- through the Recovery, Transformation and Resilience Plan NextGeneration EU within the framework of the Digital Spain 2026 Agenda.

### v2

@@ Abstract
Distillability.   Besides detecting entanglement, it is a major problem to detect whether a given bipartite state can be distilled to a two-qubit Bell pair by using multiple copies [45, 48]. Here we address this problem as follows: in Proposition 1 consider the case $S^{1}=A$ and $S^{2}=B$, both with $k$ parties having the same dimensions $d_{A}$ and $d_{B}$, similarly as in the multi-copy scenario above. Let us set $R^{(1)}=P_{A}\Pi^{\otimes k}$ and $R^{(2)}=Q_{B}\Pi^{\otimes k}$, where $P_{A}$ and $Q_{B}$ are projectors onto subspaces of ${(\mathds{C}^{d_{A}})}^{\otimes k}$ and ${(\mathds{C}^{d_{B}})}^{\otimes k}$, and $\Pi$ is a local projector onto $\mathds{C}^{2}$ (see Tab. I and Fig. 2 in [22] for specific results derived in this case). Then the (k1)(k-1)-copy action of the multilinear positive maps $\Psi_{A_{1}...A_{k-1}}$ and $\Psi_{B_{1}...B_{k-1}}$ defined by the multipartite Choi matrices $R^{(1)}$ and $R^{(2)}$, namely $\Psi_{A_{1}...A_{k-1}}\otimes\Psi_{B_{1}...B_{k-1}}(\varrho_{AB}^{\otimes k-1})$, projects k1k-1 copies of a bipartite state to the two-qubit subspace $\mathds{C}^{2}\otimes\mathds{C}^{2}$, where the PPT-criterion detects both entanglement and distillability [49, 45].
Thus the framework introduced in this work naturally provides a criterion for (k1)(k-1)-copy distillability, since two-qubit entanglement can be distilled from a state $\varrho_{AB}$ if and only if it can be transformed into a two-qubit entangled state through local projections on (k1)(k-1) copies [50, 51]. To assure that this is the case, it is sufficient to find local projections on (k1)(k-1) copies whose output is a state with non-positive partial transposition. In particular, this is guaranteed if
$\min_{\tau}\operatorname{tr}\Big((P_{A}\cdot\Pi_{A_{1}}^{\otimes k})\otimes(P_{B}\cdot\Pi_{B_{1}}^{\otimes k})\cdot\varrho_{A_{1}B_{1}}^{\otimes k-1}\otimes\tau_{A_{k}B_{k}}\Big)<0.$ (10)
Here the minimization is taken over two-qubit unit trace Hermitian operators $\tau$ with positive partial transpose $\tau^{T_{A}}\geq 0$, to optimize over all two-qubit witnesses. If strict negativity in Eq. (10) is satisfied, then the projected output has non-positive partial transposition and therefore $\varrho_{AB}$ is (k1)(k-1)-copy distillable. Here the notation $P_{A}$ and $P_{B}$ highlights that these operators have support across the $k$ copies of $A_{1}$ and $B_{1}$.
This shall be illustrated in specific examples involving a single copy, considering the two-qudit antisymmetrizer $P_{1,1}^{(d)}=(\eta_{d}(\operatorname{id})-\eta_{d}(12))/2$ and a two-dimensional local projector $\Pi=\operatorname{diag}(1,1,0,...,0)$ diagonal in the computational basis: in [22] we demonstrate that most generic states in a random sample are detected to be one-copy distillable by using $R^{(1)}=P_{1,1}\Pi^{\otimes 2}$ (see [22] for a detailed numerical analysis), effectively detecting the qubit-qudit distillable entanglement present in a subspace of a two-qudit system. Moreover, choosing $R^{(1)}=R^{(2)}=P_{1,1}\Pi^{\otimes 2}$, the operator $\mathcal{W}=R^{(1)}_{A_{1}A_{2}}\otimes R^{(2)}_{B_{1}B_{2}}$ detects both entanglement and one-copy distillability on all Werner states $\varrho_{W}$ known to be distillable, namely
@@
within $p\in(p_{0},1/2]$ with $p_{0}=\frac{d+1}{4d-2}$ [48], where $P_{2,0}^{(d)}=(\eta_{d}(\operatorname{id})+\eta_{d}(12))/2$, $d_{s}=d(d+1)/2$ and $d_{a}=d(d-1)/2$. Indeed, one verifies that
$\operatorname{tr}_{A_{1}B_{1}}\Big(\mathcal{W}\cdot{\varrho_{W}}_{A_{1}B_{1}}\otimes\mathds{1}_{A_{2}B_{2}}\Big)=pP_{2,0}^{(2)}/d_{s}+(1-p)P_{1,1}^{(2)}/d_{a}\,$, (12)
where $P_{2,0}^{(2)}$ and $P_{1,1}^{(2)}$ are the two-qubit symmetrizer and antisymmetrizer (projectors onto the triplet and singlet subspaces) respectively. Therefore, choosing the two-qubit SWAP operator $\tau=2|\phi^{+}\rangle\langle\phi^{+}|^{T_{A}}=\eta_{2}(12)$, we have $\operatorname{tr}(\mathcal{W}_{\tau}\varrho_{W})<0$ for $p\in(p_{0},1/2]$. This successful ansatz shows that the presented method to detect entanglement is also well suited to detect distillability, by using local projectors and optimization over two-qubit witnesses. This approach recovers a similar spirit as earlier criteria using witnesses [52] or SDP optimization [53, 54, 55, 56].
@@
Conclusions.   We have developed a technique to design new linear and nonlinear entanglement witnesses, and demonstrated its capabilities with examples. As a proof of concept we have shown that limited shared resources allow to employ witnesses unable to detect locally bound entangled states to construct new larger witnesses able to do so; and that available examples of witnesses for few-body bound entangled states can be used to construct new witnesses detecting bound entanglement in larger systems. The method applies to construct nonlinear witnesses acting on multiple copies. We are able to optimize these to target states with little computational resources, which is a current challenge. As a case study of both theoretical and experimental interest, we have seen that our previous multicopy construction provided in [20] can be tuned to significantly better detect noisy entangled states, while the desired properties of unitary invariance and implementability with randomized measurements have been shown to be preserved. Moreover, the technique introduced leads to practical recipes to detect distillability, which are demonstrated to be effective for relevant families of states.
@@
Acknowledgements:   The author is thankful to Paweł Horodecki, Felix Huber, Robin Krebs, Ferran Riera-Sàbat, Some Sankar Bhattacharya and Anna Sanpera for comments and support. An anonymous Referee is acknowledged for asking whether entanglement structures beyond full-separability can be detected. Support by the Foundation for Polish Science through TEAM-NET project POIR.04.04.00-00-17C1/18-00 and by NCN QuantERA Project No. 2021/03/Y/ST2/00193 is gratefully acknowledged. The author also acknowledges financial support from Spanish MICIN (projects: PID2022:141283NBI00;139099NBI00) with the support of FEDER funds, the Spanish Goverment with funding from European Union NextGenerationEU (PRTR-C17.I1), the Generalitat de Catalunya, the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project -Quantum Spain Project- through the Recovery, Transformation and Resilience Plan NextGeneration EU within the framework of the Digital Spain 2026 Agenda.
Appendix A Examples of Young projectors
The three-qudit Hilbert space $\mathcal{H}=\mathds{C}^{d}\otimes\mathds{C}^{d}\otimes\mathds{C}^{d}$ decomposes into the symmetric, standard, and antisymmetric irreducible subspaces. These are labeled by the partitions $(3,0,0)$, $(2,1,0)$ and $(1,1,1)$, and are defined by the following Young projectors,
$P_{3}$ $=\frac{1}{6}\eta_{d}\Big(\big[(\operatorname{id})+(12)+(13)+(23)+(123)+(132)\big]\Big)\,$, (13)
$P_{2,1}$ $=\frac{1}{6}\eta_{d}\Big(\big[2(\operatorname{id})-(123)-(132)\big]\Big)\,$, (14)
$P_{1^{3}}$ $=\frac{1}{6}\eta_{d}\Big(\big[(\operatorname{id})-(12)-(13)-(23)+(123)+(132)\big]\Big)\,.$ (15)
Here permutations are written in cyclic notation. For example,
$\eta_{d}(123)\mathinner{|v_{1}\rangle}\mathinner{|v_{2}\rangle}\mathinner{|v_{3}\rangle}=\mathinner{|v_{3}\rangle}\mathinner{|v_{1}\rangle}\mathinner{|v_{2}\rangle}\,.$ (16)
Appendix B New witnesses for MPPT states
Here we provide a simple example of an indecomposable three-partite entanglement witness lifted to a new entanglement witnesses detecting MPPT states in larger systems. We start from the three-qubit indecomposable witness
$W_{I}=\begin{pmatrix}\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&1\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&1&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&-1&\cdot&\cdot\\ \cdot&\cdot&\cdot&a&1&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&1&b&\cdot&\cdot&\cdot\\ \cdot&\cdot&-1&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&1&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ 1&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\end{pmatrix}$ (17)
with .=0.=0 and $ab=8$ given in [40], which is able to detect three-qubit states with local positive partial transpositions. Notice that for four-partite states this condition is not enough for MPPT, as it does not ensure positivity of the partial transposition in subsets with more than one party. The method introduced here allows to do so. Using Eq. (5) with $\tau=\phi^{+}$ and $M=|GHZ_{2,l}\rangle\langle GHZ_{2,l}|$, for $l=3$ and $l=4$ we obtain respectively a 4-qubit and 5-qubit witness whose minimial expectation value over the set of MPPT states satisfies $\operatorname{tr}(\mathcal{W}\varrho)<0$. By setting $M$ to be a $W$-state [57] or a Haar random state, a negative minimal expectation value is also obtained over the set of MPPT states.